# PartCraft: Crafting Creative Objects by Parts

### v1

@@ Abstract
This paper propels creative control in generative visual AI by allowing users to “select”. Departing from traditional text or sketch-based methods, we for the first time allow users to choose visual concepts by parts for their creative endeavors. The outcome is fine-grained generation that precisely captures selected visual concepts, ensuring a holistically faithful and plausible result. To achieve this, we first parse objects into parts through unsupervised feature clustering. Then, we encode parts into text tokens and introduce an entropy-based normalized attention loss that operates on them. This loss design enables our model to learn generic prior topology knowledge about object’s part composition, and further generalize to novel part compositions to ensure the generation looks holistically faithful. Lastly, we employ a bottleneck encoder to project the part tokens. This not only enhances fidelity but also accelerates learning, by leveraging shared knowledge and facilitating information exchange among instances. Visual results in the paper and supplementary material showcase the compelling power of PartCraft: Crafting Creative Objects by Parts in crafting highly customized, innovative creations, exemplified by the “charming” and creative birds in Fig. 1. Code is released at https://github.com/kamwoh/partcraft.
@@ 1 Introduction
Figure 1: People often form creative concepts based on existing ones [67, 44, 56, 10]. For instance, a bird enthusiast may want to craft a unique bird with different parts (e.g., heads, bodies, and wings) from common bird types (e.g., blue jay, cardinal and song sparrow). (a) Using text prompts in T2I models often results in a lack of control and deviation from the intended details, especially those visual details that are difficult to describe. (b) While sketching is a direct way, not everyone possesses the ability to sketch, particularly in intricate detail. (c) Even with reference images, existing methods (e.g., DreamBooth [55]) did not consider learning object parts, thus cannot generate with desired parts. (d) Using additional control is even cumbersome, requiring too many inputs! (e) We aim to create an object by simply selecting desired parts. PartCraft: Crafting Creative Objects by Parts learns from visual examples to generate the object with a faithful holistic structure, seamlessly integrating the chosen parts into a natural and coherent entity.
Humans are creators; AI, on the other hand, hallucinates. Creativity, arguably, is the very force driving humanity forward. Recently, generative AI has garnered considerable attention for its perceived “creativity” [54, 51, 7, 21, 45, 57, 20, 49]. Despite its promise, the challenge of control has swiftly surfaced – how can humans infuse their creativity into the generation process and regulate the extent to which AI hallucinates?
@@
Our contributions are as follows: (i) We pioneer a unique approach for fine-grained part-level control in Text-to-Image (T2I) models, empowering users to craft objects by selecting desired parts. This method streamlines the creative process, marking a significant advancement in our capacity to manipulate and reimagine visual content. (ii) We introduce PartCraft: Crafting Creative Objects by Parts, a technique that autonomously parses object parts and orchestrates them from different visual concepts, resulting in the faithful creation of a novel object. (iii) For enhanced part disentanglement and generation fidelity, we propose an entropy-based normalized attention loss and leverage a bottleneck encoder. (iv) We present two problem-specific quantitative metrics. Comprehensive experiments on CUB-200-2011 (birds) and the Stanford Dogs dataset demonstrate the superior performance of our method in generating novel objects, surpassing alternative approaches in both qualitative and quantitative evaluations.
@@ 2 Related Work
Text-to-image generation. Recent advancements in large text-to-image (T2I) diffusion models [54, 51, 7, 21, 45, 57, 20] have made significant improvements over conventional methods [74, 61, 69, 72, 48, 38] in producing high-fidelity images from text prompts. Their application scope has expanded to include both global [11, 43, 33] and localized [18, 70, 29, 5] image editing tasks, demonstrating versatility. Methods such as [25, 6, 39, 73, 8, 35, 14, 71, 24, 16, 68] have introduced more granular spatial control, such as incorporating semantic segmentation masks or bounding boxes, into large pretrained diffusion models to guide image generation. Contrary to these approaches, we focus on enhancing the control by enabling a straightforward discrete selection of desired parts. We minimize the complexity and manual intervention required by spatial controls, yet the model can compose selected parts as a coherent object seamlessly. This not only simplifies the user’s role in the generative process but also ensures that the compositional logic and coherence are inherently managed by the model’s ability. As such, our research is distinguished by focusing on the model’s inherent ability to understand and apply part relationships instead of providing additional controls.
@@ 3 Methodology
Figure 3: Overview of our PartCraft: Crafting Creative Objects by Parts. All parts are organized into a dictionary, and their semantic embeddings are learned through a textual inversion approach. For instance, a text description like “a photo of a [Head,42] [Wing,87]…” guides the optimization of the corresponding textual embedding by reconstructing the associated image. To improve generation fidelity, we incorporate a bottleneck encoder $f$ (MLP) to compute the embedding $y$ (Eq. (4)) as input to the text encoder. To promote disentanglement among learned parts, we minimize a specially designed attention loss, denoted as $\mathcal{L}_{attn}$.
Given a set of unlabeled images depicting the same object (e.g., bird) with different part details, we aim to train a T2I generative model that decomposes parts of objects into text tokens and can recompose them in a novel way. To that end, we propose PartCraft: Crafting Creative Objects by Parts, as depicted in Fig. 3.
@@ 4.1 Part Reconstruction and Image Quality Evaluation
Method Birds: CUB-200-2011 Dogs: Stanford Dogs FID CLIP DINO EMR CoSim FID CLIP DINO EMR CoSim Textual Inversion [26] 10.10 0.784 0.607 0.305 0.842 23.36 0.652 0.532 0.218 0.754 DreamBooth [55] 12.94 0.775 0.594 0.355 0.856 22.65 0.660 0.563 0.275 0.777 Custom Diffusion [37] 37.61 0.694 0.504 0.338 0.833 42.41 0.593 0.491 0.253 0.755 Break-a-Scene [4] 20.05 0.742 0.549 0.390 0.854 24.20 0.633 0.532 0.300 0.775 PartCraft: Crafting Creative Objects by Parts (Ours) 12.86 0.783 0.618 0.460 0.882 16.92 0.669 0.573 0.358 0.796
@@
(e) PartCraft: Crafting Creative Objects by Parts
@@
In Tab. 2, we summarize the performance of respective methods on the bird and dog generation, respectively. We highlight four observations: (i) Textual Inversion performs quite well compared to DreamBooth, CustomDiffusion, and Break-a-scene in terms of FID, CLIP, and DINO scores although did not fine-tune the diffusion model $\epsilon_{\theta}$. This may be due to the potential risk of overfitting when fine-tuning $\epsilon_{\theta}$ especially when learning a vast array of new concepts with many update iterations. It is also not uncommon to carefully tune the learning rate and the training iterations in these models when fine-tuning new concepts (e.g., only 800-1000 steps of updates to learn a new concept in [4]). (ii) Nonetheless, fine-tuning the diffusion model $\epsilon_{\theta}$ can help improve the ability to follow prompts as shown by increased EMR and CoSim scores (e.g., EMR of at least 5% in DreamBooth). (iii) Break-a-scene has a better ability to reconstruct the parts as shown by EMR and CoSim, this is due to the attention loss explicitly forcing the parts to focus on the respective semantic region. (iv) PartCraft: Crafting Creative Objects by Parts achieves the best performance in DINO, EMR, and CoSim scores (e.g., 7% better in EMR compared to Break-a-scene). This indicates that not only does our image-generation ability perform comparably well with textual inversion, but PartCraft: Crafting Creative Objects by Parts is also able to disentangle the parts learning so that it can follow the prompt instructions more accurately to generate the parts in a cohort. In Fig. 5, we present generated images from different methods, with CustomDiffusion exhibiting high-contrast images, possibly due to unconstrained fine-tuning on the cross-attention components $K/V$ and resulting in worse FID scores.
@@ 4.2 Part Composition Evaluation
In this section, we assess the part composition ability of different methods. In this experiment, we generate the image by mixing different parts from different species. Our findings, as shown in Fig. 6, can be summarized as follows: (i) As the number of composited parts increases, EMR and CoSim decrease, reflecting the challenge of composing multiple diverse parts. (ii) Break-a-scene and PartCraft: Crafting Creative Objects by Parts achieve notably higher EMR and CoSim scores, thanks to disentanglement through attention loss minimization. (iii) PartCraft: Crafting Creative Objects by Parts outperforms Break-a-scene significantly by token bottleneck and tailored attention loss.
@@ 4.3 Ablation Studies
(c) PartCraft: Crafting Creative Objects by Parts (Ours)
@@
Part word embedding space. We visualize the word embeddings of learned tokens of Textual Inversion [26], Break-a-scene [4] and our PartCraft: Crafting Creative Objects by Parts for birds generation (CUB-200-2011 [64]) through tSNE [41] in Fig. 11. In our PartCraft: Crafting Creative Objects by Parts, the word embeddings are the projected embeddings through Eq. (4). We can see that our projected version has a better semantic meaning such that the part embeddings are clustered together by their semantic meaning (e.g., head). We believe this is one of the reasons that our PartCraft: Crafting Creative Objects by Parts outperforms previous methods in which we can compose all parts seamlessly yet with higher quality.
Convergence analysis. We present a visual comparison of images generated by various methods in Fig. 10, spanning from the initial to the final stages of training. Notably, our PartCraft: Crafting Creative Objects by Parts demonstrates an ability to learn new concepts at even the early stages of training. Without the bottleneck encoder, we observe that the learning speed drops significantly (as evidenced by generating wrong part detail), indicating how the bottleneck serves as an important component when learning new tokens that have shared properties (e.g., common part).
@@
Transferability for Creativity. (i) In Fig. 12, we demonstrate that not only it can compose parts within the training domain (e.g., birds), but it can also transfer the learned parts to and combine with other domains (e.g., cat). This enables the creation of unique combinations, such as a cat with a dog’s ear. (ii) Leveraging the prior knowledge embedded in Stable Diffusion, PartCraft: Crafting Creative Objects by Parts can also repurpose learned parts for creative image generation. An example of this is the generation of a bird-shaped robot adorned with various parts. These examples showcase PartCraft: Crafting Creative Objects by Parts’s immense potential for diverse and limitless creative applications. Please see the supplementary material for more examples.
@@ 5 Conclusion
We propose a new way of control in generative AI. Instead of text or sketch, we “select” desired parts to create an object. We addressed the challenge of learning parts in T2I models by introducing a customized attention loss. This loss serves a dual purpose: to ensure parts are at the right location and to ensure each location is occupied by not more than one part. This greatly improves the part disentanglement. We further employ a non-linear bottleneck encoder to improve generation fidelity. Our model, PartCraft: Crafting Creative Objects by Parts, can seamlessly compose different parts from different objects, creating objects that do not exist yet holistically correct and plausible objects by mixing them. Extensive experiments demonstrated PartCraft: Crafting Creative Objects by Parts’s superior performance in both qualitative and quantitative evaluation. Moreover, the learned parts demonstrate strong transferability. We hope that our PartCraft: Crafting Creative Objects by Parts will empower artists, designers, and enthusiasts to bring the creations of their dreams to reality.
@@ 6 Limitations and Future Works
It is worth noting that the accuracy of obtained parts may be affected by using a self-supervised pre-trained feature extractor. Future work may explore the incorporation of encoders, such as [66], to improve part accuracy. We also observed challenges in composing relatively small parts, like tails and legs, which require further investigation. Additionally, we are also exploring cross-domain generation, i.e., combining learned parts from different datasets to create objects with even more diverse parts. For instance, we can merge non-rigid parts (e.g., parts from quadrupled animals) with rigid parts (e.g., parts from cars/airplanes) and form a creative structure (e.g., a car that has horse legs instead of wheels). This not only further improves the applicability of PartCraft: Crafting Creative Objects by Parts but also serves as a stepping stone to creative generation, as generative AI progresses, continually expanding the limits of achievable creativity and artistic expression.

### v2

@@ Abstract
This paper propels creative control in generative visual AI by allowing users to “select”. Departing from traditional text or sketch-based methods, we for the first time allow users to choose visual concepts by parts for their creative endeavors. The outcome is fine-grained generation that precisely captures selected visual concepts, ensuring a holistically faithful and plausible result. To achieve this, we first parse objects into parts through unsupervised feature clustering. Then, we encode parts into text tokens and introduce an entropy-based normalized attention loss that operates on them. This loss design enables our model to learn generic prior topology knowledge about object’s part composition, and further generalize to novel part compositions to ensure the generation looks holistically faithful. Lastly, we employ a bottleneck encoder to project the part tokens. This not only enhances fidelity but also accelerates learning, by leveraging shared knowledge and facilitating information exchange among instances. Visual results in the paper and supplementary material showcase the compelling power of PartCraft in crafting highly customized, innovative creations, exemplified by the “charming” and creative birds in Fig. 1. Code is released at https://github.com/kamwoh/partcraft.
@@ 1 Introduction
Figure 1: People often form creative concepts based on existing ones [70, 46, 59, 10]. For instance, a bird enthusiast may want to craft a unique bird with different parts (e.g., heads, bodies, and wings) from common bird types (e.g., blue jay, cardinal and song sparrow). (a) Using text prompts in T2I models often results in a lack of control and deviation from the intended details, especially those visual details that are difficult to describe. (b) While sketching is a direct way, not everyone possesses the ability to sketch, particularly in intricate detail. (c) Even with reference images, existing methods (e.g., DreamBooth [58]) did not consider learning object parts, thus cannot generate with desired parts. (d) Using additional control is even cumbersome, requiring too many inputs! (e) We aim to create an object by simply selecting desired parts. PartCraft learns from visual examples to generate the object with a faithful holistic structure, seamlessly integrating the chosen parts into a natural and coherent entity.
Humans are creators; AI, on the other hand, hallucinates. Creativity, arguably, is the very force driving humanity forward. Recently, generative AI has garnered considerable attention for its perceived “creativity” [57, 54, 7, 21, 47, 60, 20, 51, 52, 29]. Despite its promise, the challenge of control has swiftly surfaced – how can humans infuse their creativity into the generation process and regulate the extent to which AI hallucinates?
@@
Our contributions are as follows: (i) We pioneer a unique approach for fine-grained part-level control in Text-to-Image (T2I) models, empowering users to craft objects by selecting desired parts. This method streamlines the creative process, marking a significant advancement in our capacity to manipulate and reimagine visual content. (ii) We introduce PartCraft, a technique that autonomously parses object parts and orchestrates them from different visual concepts, resulting in the faithful creation of a novel object. (iii) For enhanced part disentanglement and generation fidelity, we propose an entropy-based normalized attention loss and leverage a bottleneck encoder. (iv) We present two problem-specific quantitative metrics. Comprehensive experiments on CUB-200-2011 (birds) and the Stanford Dogs dataset demonstrate the superior performance of our method in generating novel objects, surpassing alternative approaches in both qualitative and quantitative evaluations.
@@ 2 Related Work
Text-to-image generation. Recent advancements in large text-to-image (T2I) diffusion models [57, 54, 7, 21, 47, 60, 20, 22] have made significant improvements over conventional methods [77, 64, 72, 75, 50, 40] in producing high-fidelity images from text prompts. Their application scope has expanded to include both global [11, 45, 35] and localized [18, 73, 31, 5] image editing tasks, demonstrating versatility. Methods such as [26, 6, 41, 76, 8, 37, 14, 74, 25, 16, 71] have introduced more granular spatial control, such as incorporating semantic segmentation masks or bounding boxes, into large pretrained diffusion models to guide image generation. Contrary to these approaches, we focus on enhancing the control by enabling a straightforward discrete selection of desired parts. We minimize the complexity and manual intervention required by spatial controls, yet the model can compose selected parts as a coherent object seamlessly. This not only simplifies the user’s role in the generative process but also ensures that the compositional logic and coherence are inherently managed by the model’s ability. As such, our research is distinguished by focusing on the model’s inherent ability to understand and apply part relationships instead of providing additional controls.
@@ 3 Methodology
Figure 3: Overview of our PartCraft. All parts are organized into a dictionary, and their semantic embeddings are learned through a textual inversion approach. For instance, a text description like “a photo of a [Head,42] [Wing,87]…” guides the optimization of the corresponding textual embedding by reconstructing the associated image. To improve generation fidelity, we incorporate a bottleneck encoder $f$ (MLP) to compute the embedding $y$ (Eq. (4)) as input to the text encoder. To promote disentanglement among learned parts, we minimize a specially designed attention loss, denoted as $\mathcal{L}_{attn}$.
Given a set of unlabeled images depicting the same object (e.g., bird) with different part details, we aim to train a T2I generative model that decomposes parts of objects into text tokens and can recompose them in a novel way. To that end, we propose PartCraft, as depicted in Fig. 3.
@@ 4.1 Part Reconstruction and Image Quality Evaluation
Method Birds: CUB-200-2011 Dogs: Stanford Dogs FID CLIP DINO EMR CoSim FID CLIP DINO EMR CoSim Textual Inversion [27] 10.10 0.784 0.607 0.305 0.842 23.36 0.652 0.532 0.218 0.754 DreamBooth [58] 12.94 0.775 0.594 0.355 0.856 22.65 0.660 0.563 0.275 0.777 Custom Diffusion [39] 37.61 0.694 0.504 0.338 0.833 42.41 0.593 0.491 0.253 0.755 Break-a-Scene [4] 20.05 0.742 0.549 0.390 0.854 24.20 0.633 0.532 0.300 0.775 PartCraft (Ours) 12.86 0.783 0.618 0.460 0.882 16.92 0.669 0.573 0.358 0.796
@@
(e) PartCraft
@@
In Tab. 2, we summarize the performance of respective methods on the bird and dog generation, respectively. We highlight four observations: (i) Textual Inversion performs quite well compared to DreamBooth, CustomDiffusion, and Break-a-scene in terms of FID, CLIP, and DINO scores although did not fine-tune the diffusion model $\epsilon_{\theta}$. This may be due to the potential risk of overfitting when fine-tuning $\epsilon_{\theta}$ especially when learning a vast array of new concepts with many update iterations. It is also not uncommon to carefully tune the learning rate and the training iterations in these models when fine-tuning new concepts (e.g., only 800-1000 steps of updates to learn a new concept in [4]). (ii) Nonetheless, fine-tuning the diffusion model $\epsilon_{\theta}$ can help improve the ability to follow prompts as shown by increased EMR and CoSim scores (e.g., EMR of at least 5% in DreamBooth). (iii) Break-a-scene has a better ability to reconstruct the parts as shown by EMR and CoSim, this is due to the attention loss explicitly forcing the parts to focus on the respective semantic region. (iv) PartCraft achieves the best performance in DINO, EMR, and CoSim scores (e.g., 7% better in EMR compared to Break-a-scene). This indicates that not only does our image-generation ability perform comparably well with textual inversion, but PartCraft is also able to disentangle the parts learning so that it can follow the prompt instructions more accurately to generate the parts in a cohort. In Fig. 5, we present generated images from different methods, with CustomDiffusion exhibiting high-contrast images, possibly due to unconstrained fine-tuning on the cross-attention components $K/V$ and resulting in worse FID scores.
@@ 4.2 Part Composition Evaluation
In this section, we assess the part composition ability of different methods. In this experiment, we generate the image by mixing different parts from different species. Our findings, as shown in Fig. 6, can be summarized as follows: (i) As the number of composited parts increases, EMR and CoSim decrease, reflecting the challenge of composing multiple diverse parts. (ii) Break-a-scene and PartCraft achieve notably higher EMR and CoSim scores, thanks to disentanglement through attention loss minimization. (iii) PartCraft outperforms Break-a-scene significantly by token bottleneck and tailored attention loss.
@@ 4.3 Ablation Studies
(c) PartCraft (Ours)
@@
Part word embedding space. We visualize the word embeddings of learned tokens of Textual Inversion [27], Break-a-scene [4] and our PartCraft for birds generation (CUB-200-2011 [67]) through tSNE [43] in Fig. 11. In our PartCraft, the word embeddings are the projected embeddings through Eq. (4). We can see that our projected version has a better semantic meaning such that the part embeddings are clustered together by their semantic meaning (e.g., head). We believe this is one of the reasons that our PartCraft outperforms previous methods in which we can compose all parts seamlessly yet with higher quality.
Convergence analysis. We present a visual comparison of images generated by various methods in Fig. 10, spanning from the initial to the final stages of training. Notably, our PartCraft demonstrates an ability to learn new concepts at even the early stages of training. Without the bottleneck encoder, we observe that the learning speed drops significantly (as evidenced by generating wrong part detail), indicating how the bottleneck serves as an important component when learning new tokens that have shared properties (e.g., common part).
@@
Transferability for Creativity. (i) In Fig. 12, we demonstrate that not only it can compose parts within the training domain (e.g., birds), but it can also transfer the learned parts to and combine with other domains (e.g., cat). This enables the creation of unique combinations, such as a cat with a dog’s ear. (ii) Leveraging the prior knowledge embedded in Stable Diffusion, PartCraft can also repurpose learned parts for creative image generation. An example of this is the generation of a bird-shaped robot adorned with various parts. These examples showcase PartCraft’s immense potential for diverse and limitless creative applications. Please see the supplementary material for more examples.
@@ 5 Conclusion
We propose a new way of control in generative AI. Instead of text or sketch, we “select” desired parts to create an object. We addressed the challenge of learning parts in T2I models by introducing a customized attention loss. This loss serves a dual purpose: to ensure parts are at the right location and to ensure each location is occupied by not more than one part. This greatly improves the part disentanglement. We further employ a non-linear bottleneck encoder to improve generation fidelity. Our model, PartCraft, can seamlessly compose different parts from different objects, creating objects that do not exist yet holistically correct and plausible objects by mixing them. Extensive experiments demonstrated PartCraft’s superior performance in both qualitative and quantitative evaluation. Moreover, the learned parts demonstrate strong transferability. We hope that our PartCraft will empower artists, designers, and enthusiasts to bring the creations of their dreams to reality.
@@ 6 Limitations and Future Works
It is worth noting that the accuracy of obtained parts may be affected by using a self-supervised pre-trained feature extractor. Future work may explore the incorporation of encoders, such as [69], to improve part accuracy. We also observed challenges in composing relatively small parts, like tails and legs, which require further investigation. Additionally, we are also exploring cross-domain generation, i.e., combining learned parts from different datasets to create objects with even more diverse parts. For instance, we can merge non-rigid parts (e.g., parts from quadrupled animals) with rigid parts (e.g., parts from cars/airplanes) and form a creative structure (e.g., a car that has horse legs instead of wheels). This not only further improves the applicability of PartCraft but also serves as a stepping stone to creative generation, as generative AI progresses, continually expanding the limits of achievable creativity and artistic expression.